OpenStudy (anonymous):

Write the point-slope form of an equation of the line that passes through (-2, 5) and has a slope of zero

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

y-y1 = m(x-x1) y-5 = 0(x-(-2)) this is answer if you need to show the substitution y-5 = 0 y=5 this is the answer if you are supposed to give the answer in slope-intercept form